OpenStudy (anonymous):

A closed container has 3.06 • 1023 atoms of a gas. Each atom of the gas weighs 1.67 • 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? (1 point)

OpenStudy (anonymous):

0.47 grams, because (3.06 + 1.67) • (1023 • 10–24) = 4.73 • 10–1 0.51 grams, because (3.06 • 1.67) • (1023 • 10-24) = 5.1102 • 10-1 4.73 grams, because (3.06 + 1.67) • (1023 • 10–24) = 4.73 5.11 grams, because (3.06 • 1.67) • (1023 • 10–24) = 5.1102

OpenStudy (anonymous):

@KamiBug